OpenStudy (yrelhan4):

four tickets marked 00, 01,10, 11 respectively are placed in a bag. A ticket is drawn at random five times, being replaced each time . the probability that the sum of the numbers on tickets thus drawn is 23?

OpenStudy (shubhamsrg):

23 = 10 + 11 + 1 + 1 + 0 = 10 + 10 + 1 + 1 + 1 = 11 + 11 + 1 + 0 + 0 are any other combos possible ?

OpenStudy (yrelhan4):

wellthe answer is 25/256. and total number of ways are 1024. so there are 100 favourable ways. :P

OpenStudy (shubhamsrg):

I haven't given the ans yet, am just saying 23 can be produced in these 3 ways right,?

OpenStudy (shubhamsrg):

the ans would be (1/4)^5 *5!/2! ( 1 + 1/3! + 1/2!)

OpenStudy (shubhamsrg):

=25/256 B|

OpenStudy (yrelhan4):

oh. yeah those are possible combos. and how did you get that?

OpenStudy (shubhamsrg):

see there are 5 draws, for each draw, probab for any number is same i.e. 1/4

OpenStudy (shubhamsrg):

1st combo = 10,11,1,1,0 we can get these in 5! /2! ways next =10,10,1,1,1 => 5!/2! 3! last = 11,11,1,0,0 => 5!/2! 2! net ptobab = (1/4)^5 * 5! /2! + (1/4)^5 * 5!/2!3! + (1/4)^5 * 5!/2! 2!

OpenStudy (yrelhan4):

hmm. got it. thank you.

OpenStudy (shubhamsrg):

glad to help.